A parallelogram has sides of length 8cm and 12cm. Given the one of its angles is 65 degrees, find the lengths of its diagonals.
Answer by Edwin McCravy(20077)


Let the length of the shorter diagonal be x:



This is a case of side-angle-side given:
Use the law of cosines on the triangle with sides 8,12, and x

x² = 8² + 12² - 2*8*12*cos(65°)
x² = 64 + 144 - 192*cos(65°)
x² = 208 - 192(.4226182617)
x² = 126.8572937
x = 11.26309432

Next we use the fact that two adjacent angles of a parallelogram
are supplementary to find that the angle on the lower right is
180° - 65° or 115°. We draw the longer diagonal, and label its
length y:



This is also a case of side-angle-side given:
Use the law of cosines on the triangle with sides 8,12, and y

y² = 8² + 12² - 2*8*12*cos(115°)
y² = 64 + 144 - 192*cos(115°)
y² = 208 - 192(-.4226182617)
y² = 208 + 192(.4226182617)

y² = 289.1427063
y = 17.000419672

Edwin